Please help me solve this equation for a parallel line and a perpendicular line: y=1/4x-2; where the pass through points are (8, -1)
Thank you so much in advance,
C. Strickland

Answer by macston(5194)

.
Parallel line:
Slope will be same as given line (m=1/4)
Using slope-intercept form:
y=mx+b
y=(1/4)x+b Substitute values of given point (8,-1) to get b.
-1=(1/4)8+b
-1=2+b
-3=b
ANSWER 1: Equation of the parallel line through the given point:
.
y=(1/4)x-3
.
Perpendicular line:
Slope will be negative reciprocal of first slope (m=-4/1)
Using slope-intercept form:
y=mx+b
y=-4x+b Substitute values of given point (8,-1) to get b.
-1=-4(8)+b
-1=-32+b
31=b
ANSWER 2: Equation of perpendicular line through given point:
.
y=-4x+31